OpenStudy (anonymous):

Solve for y 0.4x-0.3y+12=0 says the answer is y=4/3x+40 I did 0.4x-0.3y=-12 -0.3y=-12-0.4x y=-12-0.4x/-0.3?

OpenStudy (anonymous):

one idea might be to multiply everything by 10 and start with 4x - 3 y + 120 = 0

OpenStudy (anonymous):

you solving for y right?

OpenStudy (anonymous):

yes, why would you multiply by 10?

OpenStudy (anonymous):

to get rid of the annoying decimals you don't have to do that, you will get the same answer working with the decimals lets do it both ways

OpenStudy (anonymous):

after multiplying by 10 and getting 4x - 3 y + 120 = 0 add 3y to both sides and get 4x + 120 = 3y divide both sides by 3 and get 4/3 x + 40 = y

OpenStudy (anonymous):

ohhh gotcha!

OpenStudy (anonymous):

i understand it after multiplying by 10. the decimals threw me off lol

OpenStudy (anonymous):

working with the decimals, add .3y to both sides and get .4x + 12 = .3y divide both sides by .3 and get .4/.3 x +12/.3 = y but .4/.3 = 4/3 and 12/.3 = 120/3 = 40 so you still get 4/3 x + 40 = y

OpenStudy (anonymous):

yeah decimals are a pain, that is why it is nice when you can eliminate them but if you multiply by 10 make sure to multiply everything by 10 in this case you have to remember to multiply 12 by 10 to get 120

OpenStudy (anonymous):

can you help me with 4/5x-2/3x+8=0 I'm stuck with this one too.

OpenStudy (anonymous):

are the x's outside of the fraction?

OpenStudy (anonymous):

(4/5)x-(2/3)y+8=0

OpenStudy (anonymous):

oh that is a y!

OpenStudy (anonymous):

multiply everything by the least common multiple of 5 and 3, which is 15

OpenStudy (anonymous):

yeah sorry typo :)

OpenStudy (anonymous):

by multiply, i really means cancel what you can so if you multiply all by 15 you get 12z - 10 y + 120 = 0

OpenStudy (anonymous):

typo there, the z should be an x

OpenStudy (anonymous):

then subtract 120 to both sides?

OpenStudy (anonymous):

you are solving for y right?

OpenStudy (anonymous):

i would add 10 y to both sides and get 12x + 120 = 10y

OpenStudy (anonymous):

Yes, I got it! :) y=(6/5)x+12

OpenStudy (anonymous):

yes, that is the right one hope it is clear why and how i multiplied all by 15

OpenStudy (anonymous):

what about (y-2)/(3)=(x+3)/(4) ? Do I multiply everything by 12?

OpenStudy (anonymous):

you can do that, yes

OpenStudy (anonymous):

so, 12y-24-12x+36?

OpenStudy (anonymous):

don't forget parentheses get 4( y - 2 ) = 3( x + 3)

OpenStudy (anonymous):

12y-24=12x+36 * typo

OpenStudy (anonymous):

errr um what?^

OpenStudy (anonymous):

when you multiply by 12 you are cancelling

OpenStudy (anonymous):

hold on i will write it out

OpenStudy (anonymous):

I figured it out! :)

OpenStudy (anonymous):

be a lot easier of the stupid equation thing was working here goes (y-2)/(3)=(x+3)/(4) multiply by 12 get 12(x - 2)/3 = 12( x + 3)/4 now cancel 3 goes in to 12 4 times, and 4 goes in to 12 3 times oh good

OpenStudy (anonymous):

y=3/4x+17/4

OpenStudy (anonymous):

yay